OpenStudy (anonymous):

Two stones are thrown vertically upward from the ground, one with three times the initial speed of the other. Assume free fall. If the faster stone takes 12.0s to return to the ground, how long will it take the slower stone to return? i got the answer it 4s If the slower stone reaches a maximum height of H, how high (in terms of H) will the faster stone go? this is my Q need help

OpenStudy (atlas):

v^2 = u^2 + 2as final speed 0 at max height so, u^2 /2a = |s| when u -> initial speed becomes 3 times, the height-> s becomes 3^2 = 9 times

OpenStudy (atlas):

v = u + at again v=0 at max height u/a = |t| so if u becomes 3 times, t-> time too becomes 3 times

OpenStudy (anonymous):

so whats the answer